OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4.

OpenStudy (anonymous):

Just simplify both sides of the equation.

OpenStudy (anonymous):

Use the Distributive Property.

OpenStudy (anonymous):

So... 3x + 3 = -2x + 3 - 4

OpenStudy (anonymous):

3x + 3 = -2x - 1 Simplify it from here.

TheSmartOne (thesmartone):

3x + 3 = -2x + 2 - 4 3x+3= -2x - 2 3x+2x+3=-2x + 2x -2 5x +3 =-2 5x +3 -3 = -2-3 5x= -5 x=-1

TheSmartOne (thesmartone):

You messed up @*Shadow* in simplifying it…

OpenStudy (anonymous):

Wow.. lol I messed up that number X_X

TheSmartOne (thesmartone):

It is ok… Mistakes happen... We are humans after all